# Your Token Becomes Worthless: Unveiling Rug Pull Schemes in Crypto Token via Code-and-Transaction Fusion Analysis

### v1

@@ 3.3.1. Rug Pull Code Risk
3.3.2. Rug Pull Transaction Behavior
@@
Behavior-1: Abnormal Network Structure. When token transfers are visualized as a transaction graph, Rug Pulls often exhibit abnormal network structures. These networks often display centralized control points, where certain nodes exhibit unusually high connectivity and centrality, suggesting a dominant role in directing liquidity and token flows. To validate this observation, we constructed token transfer graphs for Rug Pull tokens and normal tokens in our dataset and analyzed several key structure indicators, detailed in Section 4.3.2. We compared the distribution of these indicators across two types of graphs and observed significant differences, indicating the abnormal network structure in Rug Pulls. For instance, Indegree Centrality and Outdegree Centrality (Wu et al., 2024) for the nodes in the Rug Pull transfer graphs are six times higher than those in normal transfer graphs. Thus, analyzing influential nodes and dense clusters can help identify potential Rug Pull activities.
@@
Behavior-2: Market Manipulation. Another prevalent behavior in Rug Pulls is market manipulation, aimed at creating a deceptive appearance of liquidity and trading activity. This manipulation typically involves addresses controlled by the scammers, working collaboratively to purchase large quantities of tokens within a short time period. These transactions create the illusion of high demand and rapid price growth, luring unsuspecting investors into buying the token at inflated prices. To further disguise the manipulation, scammers may distribute the transactions across various addresses to simulate trading activity, making it difficult to distinguish between legitimate price increases and artificial price manipulation. For example, we analyzed the trading activity of NexF (Nex, 2025), a Rug Pull token, and found a remarkable symmetry between the buy and sell volumes on the Uniswap. Specifically, over a 20-day period, many addresses purchased the token and sold it afterward, inflating the trading volume and creating an illusion of increased liquidity.
4. METHODOLOGY
4.1. Overview
@@
Figure 2 shows the overall design of RPHunter, consisting of three main modules: Semantic Risk Code Graph (SRCG) Construction, Token Flow Behavior Graph (TFBG) Construction, and Rug Pull Detection. Given a crypto token, RPHunter initially constructs the semantic code graph of the token contract, followed by establishing declarative rules for each code risk and performing flow analysis to extract code risk information. These extracted critical operations and relationships are then used to construct a semantic risk code graph. Concurrently, RPHunter filters all token flow transactions and leverages the raw data to extract the features of accounts and transactions based on abnormal transaction behaviors: abnormal network structure and market manipulation. These features are used to construct a token flow behavior graph. Finally, RPHunter leverages graph neural networks to represent features of SRCG and TFBG. An attention fusion model is employed to balance these features, enabling a comprehensive detection of Rug Pulls.
Figure 2. Overview of RPHunter
4.2. Semantic Risk Code Graph Construction
Figure 3. Relations and Declarative Rules
4.2.1. Code Preprocessing
@@
Given a crypto token address, RPHunter firstly uses the Web3 API (Web, 2025) getCode to fetch the bytecode of the contract. Then it leverages the open-source binary lifting decompiler Gigahorse (Grech et al., 2019) to decompile the contract bytecode and generate the intermediate representation (IR). The IR of the contract consists of a series of basic blocks. Each block encapsulates a sequence of statements. Furthermore, the control flow between two blocks is determined based on the predecessor and successor IDs. A basic block may have multiple predecessors and successors, indicating complex conditional relationships. Based on the IR of a contract, RPHunter employs a depth-first search (DFS) method for each function to traverse and construct the control flow graph (CFG). Simultaneously, RPHunter analyzes relationships between data definitions and usage, introducing data dependency relationships. Finally, RPHunter constructs a directed semantic code graph (SCG), denoted as $SCG=(B,E)$, where $B$ is the set of nodes representing the basic blocks and $E$ is the set of edges indicating the control flow and data flow relationships between these blocks.
4.2.2. Code Risk Information Extraction
@@
Relations. Here, we use s to denote a statement, v to denote a variable, f to denote a function, and m to denote a mapping structure. To better represent core operations and relationships within the code, we design some basic relations shown in Figure 3. The five basic relations shown in Figure 3 are induced with the assistance of Gigahorse. For instance, DF(v1, v2) captures the data-flow dependency between v1 and v2, and VC(s, v1, v2) detects whether a comparison operation is performed between v1 and v2 within s.
@@
Furthermore, according to the summarized code risks, Rug Pull code risks predominantly occur within privileged functions and token transfer functions. Thus, we design some function relations that encapsulate the properties and operations. For example, PFAV(f, v) recognizes the public function f and identifies its argument variable v the function passes. FUMV(f, s, v) identifies function f that utilizes a storage variable v within s, which can be modified by privileged addresses. As the focus of this paper does not revolve around the induction of these relations, the details of their induction calculations are not furnished in this manuscript. More information can refer to Gigahorse (Grech et al., 2019) and our prototype tool (RPhunter, 2025).
Declarative Rules. To effectively capture critical nodes and edges potentially indicative of code risk, we establish “Basic Components ($BC$) + Plugins ($Pi$)” declarative rules shown in Figure 3. Each of the three major categories of code risk is represented by a basic component, which captures the shared characteristics common across their respective subcategories. Tailed to each of the eight subcategories of code risks, specific plugins are developed to detect unique patterns specific to that subcategory, ensuring a more fine-grained analysis of the code risks.
Using Address Restrict (ADDR) as an example, the category Sale Restrict is primarily observed within token transfer function that conceals mechanisms restricting token sales. In the basic component termed as C1-BC, we first utilize relation PTF to identify token transfer function. These restricted token sales based on amount, timestamp, or address are controlled by the privileged addresses, involving storage variables and mappings. Then we extract these storage variables and mappings controlled by privileged addresses, denoted as tainted data (PUMV $\lor$ PUMD). Additionally, we perform flow propagation analysis on these storage variables and mappings to identify other tainted data. In ADDR, the plugin termed as C1-P3 treats the arguments of token transfer function (i.e., address variable) as sensitive variables (PFAV). By analyzing flow propagation, it examines whether these sensitive variables are compared against the tainted data defined in C1-BC, indicating potential constraints (VC). Notably, due to direct comparison operators like “¿” or “¡” are not used in address restriction (C1-P3), the comparison operation analysis relies on the IZ(s, v) to check whether specific conditions are compared to zero.
Flow Analysis. Based on the declarative rules outlined, we propose a flow analysis framework to identify the semantics of code risk effectively. Specifically, given a SCG of the token, we combine the basic component and plugin of each code risk, denoted as $Ci:$$BC$land $Pj$, to capture key operations and relationships. An example of Address Restrict (ADDR) is illustrated in the Figure 4.
@@
Figure 4. Example of Flow Analysis
4.2.3. Semantic Risk Code Graph Generator
@@
4.3. Token Flow Behavior Graph Construction
4.3.1. Transaction Preprocessing
Based on the empirical study in Section 3, it is noted that token flow can be an important indicator of suspicious transaction behaviors. Consequently, our primary focus is on transactions involving token flows. Given a token contract address, we leverage the blockchain explorer to collect transaction receipts involving the token flow. We focus on the earliest 500 transactions (Wu et al., 2024), as these initial transactions can help capture suspicious transaction behavior before the occurrence of Rug Pull. A detailed discussion on the impact of the number of selected transactions is provided in Section 5.2. For each transaction receipt, essential attributes such as the transfer value and timestamp are extracted as raw data to characterize the transaction. However, the variation in decimal places across different tokens can cause differences in the magnitudes of transfer values. Thus, token transfer values are processed in two steps. First, values are normalized using decimal scaling $v^{{}^{\prime}}=v/10^{dec}$, where $v^{{}^{\prime}}$ and $v$ denote the normalized and origin value. $dec$ denotes the number of decimal places in the smallest unit of the token. Second, a logarithmic transformation is applied to these values, helping mitigate the impact of extreme outliers.
4.3.2. Token Flow Behavior Graph Generator
@@
Based on the Rug Pull transaction behaviors: Abnormal Network Structure(ANS) and Market Manipulation(MM), outlined in Section 3.3.2, we extract features from two key perspectives: network structure and manipulative behavior. The network structure features are designed to capture graph properties that may indicate ANS, while the manipulative behavior features focus on identifying outliers in specific attributes that could indicate suspicious MM.
Node Feature Extraction. Node features can be extracted from both network structure and manipulative behavior two perspectives. The node features are shown in Table 2. In terms of the network structure, we delve into the structural node attributes by calculating centrality measures and clustering coefficients, including degree centrality, indegree centrality, outdegree centrality, betweenness centrality, closeness centrality, eigenvector centrality, katz centrality, and clustering coefficient. These metrics provide insights into the node prominence and connectivity within the network, helping identify potential manipulation nodes.
Table 2. Extracted Node and Edge Features in Constructing TFBG
@@
Node Structure Centrality and Clustering Capture node importance from different views
@@
Short-term Max Incoming Tx The maximum amounts and numbers of incoming Tx over a short time period
Short-term Max Outgoing Tx The maximum amounts and numbers of outgoing Tx over a short time period
@@
Investment Cumulative Incoming Tx Total amounts and numbers of incoming Tx before the edge’s timestamp
Cumulative Outcoming Tx Total amounts and numbers of outgoing Tx before the edge’s timestamp
Short-term Max Incoming Tx The maximum amounts and numbers of cumulative incoming Tx over a short time period
Short-term Max Outgoing Tx The maximum amounts and numbers of cumulative outgoing Tx over a short time period
Edge Feature Extraction. Edge features are extracted solely from the manipulative behavior perspective to characterize interactions between nodes. Edge features about manipulative behavior are classified into three categories: time-series, transaction, and investment, outlined in Table 2. Time-series features include three indicators. CreationInterval measures the time elapsed from token contract creation to the transaction. A short interval may indicate premeditated actions immediately after contract deployment. LatestInterval measures the time interval between the current transaction and the last transaction. Short intervals in high-amount trading may suggest wash trading common in Rug Pull scenarios. IfApprove checks whether an approve operation preceded a transferFrom transaction. Transfers lacking prior approval might indicate unauthorized token movements, raising concerns about potential privileged operations.
@@
4.4. Rug Pull Detection
4.4.1. Code Graph Embedding Learning
Prior work (Yu et al., 2020) has revealed the success of the Pre-train Bert model in handling program instructions. To effectively capture the semantic information, we resort to the Bert network for encoding the nodes within the SRCG. For each statement, RPHunter strips away the statement identifier and operands, retaining only the opcode. Consequently, the content of each block is distilled into a sequence of opcodes. Subsequently, RPHunter employs the Byte Tokenizer (Sennrich, 2015) for opcode tokenization and utilizes the BERT model (Devlin, 2018) consisting of 12 attention heads for basic block embedding, with a dimension of 36. Finally, the node embedding within the SRCG can be computed by the trained model. Moreover, we employ a heterogeneous graph neural network to embed the SRCG. Specifically, inspired by the work (Schlichtkrull et al., 2018), we implement a relational graph convolutional network (RGCN) to model the varied interactions between nodes and edges.
@@ 4.4.1. Code Graph Embedding Learning
4.4.2. Transaction Graph Embedding Learning
@@
We specifically introduce the aggregation process based on GCN. In GCN, node features are updated by aggregating the features of the node itself and its neighbors, similar to Equation 4.4.1. Unlike the RGCN approach, the convolution process here is more straightforward as it does not consider different edge and node types during the aggregation. For the edge features, we not only consider the inherent features of the edge but also devise a mask mechanism to aggregate features of edges that precede the current edge in time and share the same source or target nodes as follows.
(2) $M_{(u,v,t)}=\{e^{\prime}\in S\mid(e^{\prime}=(u,v^{\prime},t^{\prime})\lor e^{\prime}=(u^{\prime},v,t^{\prime}))\land(t^{\prime}<t)\}$
@@
(3) $\vec{E}^{l+1}_{e}=\sigma(W^{l}_{e}\cdot(\vec{E}^{l}_{e}\parallel\vec{N}^{l}_{u}\parallel\vec{N}^{l}_{v}\parallel Mean(\vec{E}^{l}_{M_{(u,v,t)}})))$
@@ 5. EVALUATION
5.1. Experiment Setup
Dataset. To ensure comprehensive evaluation, we constructed our ground-truth dataset based on two principles: (1) the availability of open-source code, which facilitates reviewing and analyzing the token contract, and (2) the presence of transaction history, which allows for the analysis of transaction behaviors. We collected a total of 1048 Rug Pull tokens that occurred before May 2024. After excluding 131 tokens without source code and 272 tokens without transaction history, we obtained 645 real-world Rug Pull tokens, detailed in Section 3.1. Additionally, to train our model for Rug Pull detection, we collected 1806 benign tokens from an open-source dataset (Wu et al., 2024). These tokens are real-world crypto tokens, which have already been manually inspected and labeled as benign tokens. We excluded 131 tokens from the benign tokens due to the lack of transactions. As a result, we obtained a ground-truth dataset with 645 Rug Pull tokens and 1675 benign tokens for our experiments. And to collect token flow transactions, we leveraged blockchain explorer API to get a list of Token Transfer Events by token address.
Implementation. We implemented the prototype of RPHunter in Python and Datalog. To extract code risk information, we defined declarative rules using Datalog based on Gigahorse (Grech et al., 2019). We implemented Rug Pull Detector using over 2000 lines of Python, with the graph neural network designed using Pytorch (Paszke et al., 2019). The training, validation and test sets were split in a 60%-20%:-20% ratio. To ensure the training set represents all samples, we used five-fold cross-validation and reported the average results.

### v2

@@ III-C1 Rug Pull Code Risk
III-C2 Rug Pull Transaction Behavior
@@
Behavior-1: Abnormal Network Structure. When token transfers are visualized as a transaction graph, Rug Pulls often exhibit abnormal network structures. These networks often display centralized control points, where certain nodes exhibit unusually high connectivity and centrality, suggesting a dominant role in directing liquidity and token flows. To validate this observation, we constructed token transfer graphs for Rug Pull tokens and normal tokens in our dataset and analyzed several key structure indicators, detailed in Section IV-C2. We compared the distribution of these indicators across two types of graphs and observed significant differences, indicating the abnormal network structure in Rug Pulls. For instance, Indegree Centrality and Outdegree Centrality [23] for the nodes in the Rug Pull transfer graphs are six times higher than those in normal transfer graphs. Thus, analyzing influential nodes and dense clusters can help identify potential Rug Pull activities.
@@
Behavior-2: Market Manipulation. Another prevalent behavior in Rug Pulls is market manipulation, aimed at creating a deceptive appearance of liquidity and trading activity. This manipulation typically involves addresses controlled by the scammers, working collaboratively to purchase large quantities of tokens within a short time period. These transactions create the illusion of high demand and rapid price growth, luring unsuspecting investors into buying the token at inflated prices. To further disguise the manipulation, scammers may distribute the transactions across various addresses to simulate trading activity, making it difficult to distinguish between legitimate price increases and artificial price manipulation. For example, we analyzed the trading activity of NexF [24], a Rug Pull token, and found a remarkable symmetry between the buy and sell volumes on the Uniswap. Specifically, over a 20-day period, many addresses purchased the token and sold it afterward, inflating the trading volume and creating an illusion of increased liquidity.
IV METHODOLOGY
IV-A Overview
@@
Fig. 2 shows the overall design of RPHunter, consisting of three main modules: Semantic Risk Code Graph (SRCG) Construction, Token Flow Behavior Graph (TFBG) Construction, and Rug Pull Detection. Given a crypto token, RPHunter initially constructs the semantic code graph of the token contract, followed by establishing declarative rules for each code risk and performing flow analysis to extract code risk information. These extracted critical operations and relationships are then used to construct a semantic risk code graph. Concurrently, RPHunter filters all token flow transactions and leverages the raw data to extract the features of accounts and transactions based on abnormal transaction behaviors: abnormal network structure and market manipulation. These features are used to construct a token flow behavior graph. Finally, RPHunter leverages graph neural networks to represent features of SRCG and TFBG. An attention fusion model is employed to balance these features, enabling an effective detection of Rug Pulls.
Figure 2: Overview of RPHunter
IV-B Semantic Risk Code Graph Construction
Figure 3: Relations and Declarative Rules
IV-B1 Code Preprocessing
@@
Given a crypto token address, RPHunter firstly uses the Web3 API [25] getCode to fetch the bytecode of the contract. Then it leverages the open-source binary lifting decompiler Gigahorse [26] to decompile the contract bytecode and generate the intermediate representation (IR). The IR of the contract consists of a series of basic blocks. Each block encapsulates a sequence of statements. Furthermore, the control flow between two blocks is determined based on the predecessor and successor IDs. A basic block may have multiple predecessors and successors, indicating complex conditional relationships. Based on the IR of a contract, RPHunter employs a depth-first search (DFS) method for each function to traverse and construct the control flow graph (CFG). Simultaneously, RPHunter analyzes relationships between data definitions and usage, introducing data dependency relationships. Finally, RPHunter constructs a directed semantic code graph (SCG), denoted as $SCG=(B,E)$, where $B$ is the set of nodes representing the basic blocks and $E$ is the set of edges indicating the control flow and data flow relationships between these blocks.
IV-B2 Code Risk Information Extraction
@@
Relations. Here, we use s to denote a statement, v to denote a variable, f to denote a function, and m to denote a mapping structure. To better represent core operations and relationships within the code, we design some basic relations shown in Fig. 3. The five basic relations shown in Fig. 3 are induced with the assistance of Gigahorse. For instance, DF(v1, v2) captures the data-flow dependency between v1 and v2, and VC(s, v1, v2) detects whether a comparison operation is performed between v1 and v2 within s.
@@
Furthermore, according to the summarized code risks, Rug Pull code risks predominantly occur within privileged functions and token transfer functions. Thus, we design some function relations that encapsulate the properties and operations. For example, PFAV(f, v) recognizes the public function f and identifies its argument variable v the function passes. FUMV(f, s, v) identifies function f that utilizes a storage variable v within s, which can be modified by privileged addresses. As the focus of this paper does not revolve around the induction of these relations, the details of their induction calculations are not furnished in this manuscript. More information can refer to Gigahorse [26] and our prototype tool [8].
Declarative Rules. To effectively capture critical nodes and edges potentially indicative of code risk, we establish “Basic Components ($BC$) + Plugins ($Pi$)” declarative rules shown in Fig. 3. Each of the three major categories of code risk is represented by a basic component, which captures the shared characteristics common across their respective subcategories. Tailed to the eight subcategories of code risks, specific plugins are developed to detect unique patterns specific to the subcategories, ensuring a fine-grained analysis of code risks.
Using Address Restrict (ADDR) as an example, the category Sale Restrict is primarily observed within token transfer function that conceals mechanisms restricting token sales. In the basic component termed as C1-BC, we first utilize relation PTF to identify token transfer function. These restricted token sales based on amount, timestamp, or address are controlled by the privileged addresses, involving storage variables and mappings. Then we extract these storage variables and mappings controlled by privileged addresses, denoted as tainted data (PUMV $\lor$ PUMD). Additionally, we perform flow propagation analysis on these storage variables and mappings to identify other tainted data. In ADDR, the plugin termed as C1-P3 treats the arguments of token transfer function (i.e., address variable) as sensitive variables (PFAV). By analyzing flow propagation, it examines whether these sensitive variables are compared against the tainted data defined in C1-BC, indicating potential constraints (VC). Notably, due to direct comparison operators like “$>$” or “$<$” are not used in address restriction (C1-P3), the comparison operation analysis relies on the IZ(s, v) to check whether specific conditions are compared to zero.
Flow Analysis. Based on the declarative rules outlined, we propose a flow analysis framework to identify the semantics of code risk effectively. Specifically, given a SCG of the token, we combine the basic component and plugin of each code risk, denoted as $Ci:$$BC$land $Pj$, to capture key operations and relationships. An example of Address Restrict (ADDR) is illustrated in the Fig. 4.
Figure 4: Example of Flow Analysis
@@
IV-B3 Semantic Risk Code Graph Generator
@@
IV-C Token Flow Behavior Graph Construction
IV-C1 Transaction Preprocessing
Based on the empirical study in Section III, it is noted that token flow can be an important indicator of suspicious transaction behaviors. Consequently, our primary focus is on transactions involving token flows. Given a token contract address, we leverage the blockchain explorer to collect transaction receipts involving the token flow. We focus on the earliest 500 transactions [23], as these initial transactions can help capture suspicious transaction behavior before the occurrence of Rug Pull. A detailed discussion on the impact of the number of selected transactions is provided in Section V-B. For each transaction receipt, essential attributes such as the transfer value and timestamp are extracted as raw data to characterize the transaction. However, the variation in decimal places across different tokens can cause differences in the magnitudes of transfer values. Thus, token transfer values are processed in two steps. First, values are normalized using decimal scaling $v^{{}^{\prime}}=v/10^{dec}$, where $v^{{}^{\prime}}$ and $v$ denote the normalized and origin value. $dec$ denotes the number of decimal places in the smallest unit of the token. Second, a logarithmic transformation is applied to these values, helping mitigate the impact of extreme outliers.
IV-C2 Token Flow Behavior Graph Generator
@@
TABLE II: Extracted Node and Edge Features in Constructing TFBG
@@
Node Structure Centrality and Clustering (8) Capture node importance from different views
@@
Short-term Max Incoming Tx (2) The maximum amounts and numbers of incoming Tx over a short time period
Short-term Max Outgoing Tx (2) The maximum amounts and numbers of outgoing Tx over a short time period
@@
Investment Cumulative Incoming Tx (2) Total amounts and numbers of incoming Tx before the edge’s timestamp
Cumulative Outcoming Tx (2) Total amounts and numbers of outgoing Tx before the edge’s timestamp
Short-term Max Incoming Tx (2) The maximum amounts and numbers of cumulative incoming Tx over a short time period
Short-term Max Outgoing Tx (2) The maximum amounts and numbers of cumulative outgoing Tx over a short time period
Based on the Rug Pull transaction behaviors: Abnormal Network Structure(ANS) and Market Manipulation(MM), outlined in Section III-C2, we extract features from two key perspectives: network structure and manipulative behavior. The network structure features are designed to capture graph properties that may indicate ANS, while the manipulative behavior features focus on identifying outliers in specific attributes that could indicate suspicious MM.
Node Feature Extraction. Node features can be extracted from both network structure and manipulative behavior two perspectives. The node features are shown in Table II. In terms of the network structure, we delve into the structural node attributes by calculating centrality measures and clustering coefficients, including degree centrality, indegree centrality, outdegree centrality, betweenness centrality, closeness centrality, eigenvector centrality, katz centrality, and clustering coefficient. These metrics provide insights into the node prominence and connectivity within the network, helping identify potential manipulation nodes.
Edge Feature Extraction. Edge features are extracted solely from the manipulative behavior perspective to characterize interactions between nodes. Edge features about manipulative behavior are classified into three categories: time-series, transaction, and investment, outlined in Table II. Time-series features include three indicators. CreationInterval measures the time elapsed from token contract creation to the transaction. A short interval may indicate premeditated actions immediately after contract deployment. LatestInterval measures the time interval between the current transaction and the last transaction. Short intervals in high-amount trading may suggest wash trading common in Rug Pulls. IfApprove checks whether an approve operation preceded a transferFrom transaction. Transfers lacking prior approval might indicate unauthorized token movements, raising concerns about privileged operations.
@@
IV-D Rug Pull Detection
IV-D1 Code Graph Embedding Learning
Prior work [27] has revealed the success of the Pre-train Bert model in handling program instructions. To effectively capture the semantic information, we resort to the Bert network for encoding the nodes within the SRCG. For each statement, RPHunter strips away the statement identifier and operands, retaining only the opcode. Consequently, the content of each block is distilled into a sequence of opcodes. Subsequently, RPHunter employs the Byte Tokenizer [28] for opcode tokenization and utilizes the BERT model [29] consisting of 12 attention heads for basic block embedding, with a dimension of 36. Finally, the node embedding within the SRCG can be computed by the trained model. Moreover, we employ a heterogeneous graph neural network to embed the SRCG. Specifically, inspired by the work [30], we implement a relational graph convolutional network (RGCN) to model the varied interactions between nodes and edges.
@@ IV-D1 Code Graph Embedding Learning
IV-D2 Transaction Graph Embedding Learning
@@
We specifically introduce the aggregation process based on GCN. In GCN, node features are updated by aggregating the features of the node itself and its neighbors, similar to Equation 1. Unlike the RGCN approach, the convolution process here is more straightforward as it does not consider different edge and node types during the aggregation. For the edge features, we not only consider the inherent features of the edge but also devise a mask mechanism to aggregate features of edges that precede the current edge in time and share the same source or target nodes as follows.
$M_{(u,v,t)}=\{e^{\prime}\in S\mid$ $(e^{\prime}=(u,v^{\prime},t^{\prime})$
$\lor e^{\prime}=(u^{\prime},v,t^{\prime}))\land\ (t^{\prime}<t)\}$ (2)
@@
$\vec{E}^{l+1}_{e}=\sigma(W^{l}_{e}\cdot(\vec{E}^{l}_{e}\parallel\vec{N}^{l}_{u}\parallel\vec{N}^{l}_{v}\parallel Mean(\vec{E}^{l}_{M_{(u,v,t)}})))$ (3)
@@ V EVALUATION
V-A Experiment Setup
Dataset. To ensure comprehensive evaluation, we constructed our ground-truth dataset based on two principles: (1) the availability of open-source code, which facilitates reviewing and analyzing the token contract, and (2) the presence of transaction history, which allows for the analysis of transaction behaviors. We collected a total of 1048 Rug Pull tokens that occurred before May 2024. After excluding 131 tokens without source code and 272 tokens without transaction history, we obtained 645 real-world Rug Pull tokens, detailed in Section III-A. Additionally, to train our model for Rug Pull detection, we collected 1806 benign tokens from an open-source dataset [23]. These tokens are real-world crypto tokens, which have already been manually inspected and labeled as benign tokens. We excluded 131 tokens from the benign tokens due to the lack of transactions. As a result, we obtained a ground-truth dataset with 645 Rug Pull tokens and 1675 benign tokens for our experiments. And to collect token flow transactions, we leveraged blockchain explorer API to get a list of Token Transfer Events by token address.
Implementation. We implemented the prototype of RPHunter in Python and Datalog. To extract code risk information, we defined declarative rules using Datalog based on Gigahorse [26]. We implemented Rug Pull Detector using over 2000 lines of Python, with the graph neural network designed using Pytorch [31]. The training, validation and test sets were split in a 60%-20%:-20% ratio. To ensure the training set represents all samples, we used five-fold cross-validation and reported the average results.